OpenStudy (anonymous):

Multivariable max-min problem: Find the max and min values of the function f(x,y)=x^2+y^2 when x, y belong to the line x+y=1. So, i followed the Lagrange multipliers method and ended with this single point: (1/2,1/2). How I'm supposed to determine whether this is a max a min o none of them?.

ganeshie8 (ganeshie8):

There is no way to know the type of extrema using lagrange multipliers method

ganeshie8 (ganeshie8):

you need to find the type by other means

OpenStudy (anonymous):

Not even with the Hessian matrix?

ganeshie8 (ganeshie8):

why do you want to use hessian matrix if figuring out the type of extrema is easy simply by staring at the function ?

OpenStudy (whpalmer4):

Pick a nearby point on the line \(x+y=1\) and evaluate \(f(x,y)\) there. Is it larger or smaller than \(f(1/2,1/2)\)? That should tell you what you need to know, no?

ganeshie8 (ganeshie8):

if you want to practice hessian matrix stuff, you may do it that way... otherwise, i feel using hessian matrix to figure out the type of extrema is a sledgehammer in present problem

OpenStudy (anonymous):

I know is a minima, but i'm looking for a general (analitic) method for the case when the constraint functions are not closed sets. For use in more complicated problems.

ganeshie8 (ganeshie8):

Ahh okay then hessian matrix should work nicely

ganeshie8 (ganeshie8):

\[|H| = \begin{vmatrix}0&1&1\\1&2&0\\1&0&2 \end{vmatrix} = -4\] so any critical point is a minimum

OpenStudy (amoodarya):

as a second approach , you can write y=1-x and put it in f=x^2+(1-x)^2 then you have single variable problem \[f=x^2+(1-x)^2=\\2x^2-2x+1\\f'=4x-2=0\\x=\frac{1}{2} \to y=\frac{1}{2}\]

OpenStudy (amoodarya):

and this is obviously min ,because f''=4>0